a student collected:
temp water bath: 90 C
barometric pressure: 765 torr
volume flask: 260 mL
unknown, some unvaporized liquid remained in flask. if unknown gas was dichloroethane (C2H4Cl2) and his calculated molar lass was 112 g/mol, what mass of liquid remained unvaporized?
1
Expert's answer
2019-07-01T03:55:23-0400

"PV= \frac {m}{MW}RT"


"m=\frac {P \cdot V \cdot MW}{R \cdot T}"


"P= \frac {760 torr}{1} \cdot \frac {1atm}{760torr} = 1atm"


"T=90+273.15=363.15 K"

"V= \frac {260}{1000}=0.260 L"


"R= 0.0821 \frac {L \cdot atm}{K \cdot mol}"


"Total \space mass= \frac {1 atm \cdot 0.260L \cdot 112 g/mol}{0.0821 \frac {L \cdot atm}{K \cdot mol} \cdot 363.15 K} = 0.977 g"


"Mass \space of \space C_2H_4Cl_2 = \frac {1 atm \cdot 0.260L \cdot 99 g/mol}{0.0821 \frac {L*atm}{K*mol} \cdot 363.15 K}=0.863 g"


"Mass \space of \space an \space unvaporized \space substance = 0.977 - 0.863 = 0.114 g"